OpenStudy (vortish):

solve the system of equations by graphing 5x + y = -8 and x +2y=2

OpenStudy (zzr0ck3r):

do you know how to frapf an equation?

OpenStudy (zzr0ck3r):

graph?

OpenStudy (vortish):

im having issues with that part of this equation

OpenStudy (zzr0ck3r):

google how to graph an equation, it will be much faster than someone trying to show you on here, once you know how to graph the equations, the answer is the point where they cross

OpenStudy (anonymous):

You need two points to graph. 5x + y = -8 The best way to find the points is by isolating the y in the equation. Like this. 5x + y = -8 y= -8 -5x Now plug in an x and find the y. The two easiest points is always x=0 and x=1 So find the the y when x=0 and find the y when x=1. That will give you the two points (X,Y) Plot them and draw a line between them and that's your graph.